OpenStudy (anonymous):

can some one help me with steps to solve this problem? {5/(a+h)+4)-(5)/(a+4)}/h

OpenStudy (anonymous):

You're lacking some parentheses. Do you mind using the equation editor? Is this a difference quotient?

OpenStudy (anonymous):

try computin\g \[\frac{5}{a+h+4}-\frac{5}{a+4}\] first don't multiply out in the denominator, just in the numerator

OpenStudy (anonymous):

all this is over h as the denominatar

OpenStudy (anonymous):

yes, but my suggestion is to subtract first, divide by \(h\) last when you subtract, a miracle will occur and you will be left with a factor of \(h\) in the top, which will cancel with the \(h\) in the denominator

OpenStudy (anonymous):

the first step is this: \[\frac{5}{x+h+4}-\frac{5}{x+4}=\frac{5(x+4)-5(x+h+4)}{(x+h+4)(x+4)}\] don't multipliy out in the denominator, just leave it alone, but do the algebra in the numerator